OpenStudy (anonymous):

given the equation: y = 5x + 6 7x + 4y = 24 solve by substitution

OpenStudy (anonymous):

7x + 4(5x+6) = 24 7x + 20x + 24 = 24 27x +24 = 24 subtract 24 on both sides 27x = 0 x=0 y=5x+6 y = 5(0) + 6 y = 0 + 6 y = 6 (0, 6)

OpenStudy (anonymous):

whats x?

OpenStudy (anonymous):

0